Find b, given that a = 18.2, B = 62°, and C = 48°
Answer by Alan3354(69443)

Find b, given that a = 18.2, B = 62°, and C = 48°
===========
Angle A = 180 - (62+48) = 70 degs
--------
Use the Sine Law:
b/sin(B) = a/sin(A)